OpenStudy (anonymous):

1.85-1.03/1.03 x 100% = ???

OpenStudy (anonymous):

\[\frac{ 1.85-1.03 }{ 1.03 }(100)=\]

OpenStudy (anonymous):

0.82/1.02 (100)= [if 1.03=1.02] 0.41/0.51 (100)= 0.40/0.50 (100)= [if 0.41 = 0.40 and 0.51 = 0.50] 4/5 (100)= 4 (20)= 80%

OpenStudy (anonymous):

http://tini.cc/f6G

OpenStudy (anonymous):

did you get it @tc1979 ??